OpenStudy (anonymous):

What is the equation of the line passing through (2, -3) and parallel to the line y = -6x - 1 in standard form? a medal for first that answers

OpenStudy (anonymous):

maybe, but i don't want to give you them because, you cna base your question off of that

OpenStudy (anonymous):

jsut give the answer please

OpenStudy (anonymous):

the multipul choice are these a. 6x + y = 9 b. 6x - y = -9 c. x + 6y = 9 d. x - 6y = -9

OpenStudy (anonymous):

Oops, sorry. It's 6x + y = 9. So A.

OpenStudy (anonymous):

Do you want the explanation?

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

never mind i got it